OpenStudy (batmano3):

Write the standard form of the equation of the line with slope of 3 that passes through the point (2, - 5)

OpenStudy (styxer):

You can use the equation of line : (y-y0)=m(x-x0) slope = m = 3 x0 = 2 y0 = -5 (y+5)=3(x-2) y+5=3x-6 y=3x-11

OpenStudy (batmano3):

Oh so then I will plug it in standard form which ax+by=c

OpenStudy (styxer):

standart form is ax+by+c=0 in this case 3x-y-11=0